Shift Summary for Eve shift, 08.31.99

Jim Graham (jagraham@hep.uchicago.edu)
Wed, 01 Sep 1999 00:11:09 -0500

Shift Summary for Eve shift, 08.31.99

On Shift: JAG, UE

Hours of beam delivered at all intensities: 8.5
Hours of beam delivered at nominal intensity: 8.5
Hours of E832 data (special runs not included): 5.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: Tape changes, kfrends, replacing CAMAC
power supply on crate 23

========================================================================
Runs ended this shift:
Run 14382 - beam 832 - 24.1 M events
Run 14383 - beam 832 - 3.96 M events
Run 14384 - beam 832 - 7.61 M events
.. some other garbage debugging runs

Tapes brought to FCC this shift:
QKK363 - 368
QKS363 - 368
QKC363 - 368

QKCC16-17, QKA020-22, QKN031-32, QKSM15-16, QKNL28-30, QKLA15

========================================================================
Accesses during shift: EB, JAG, RSK to inspect, then replace, power
supply on CAMAC crate 23 (DC 2 readout).

========================================================================
Overview of shift:
Great 832 running except for power supply that went flaky.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14382 15:30 continued from day shift
17:00 run ends on full tapes
14383 17:07 start run
18:15 multiple kfrends and a pipeline reset necessary
18:30 TE quadrant of CsI drops out. Can't get it back
with pipeline resets. Stop run to do full
drt_fresh_start.
14384 18:40 Start run
18:45 Dave N. puts 2 new bottles of ArEth on east side.
20:25 Ktevana errors, kfrend
21:05 Trigger hangs on busy. Can't find it. Kfrends don't
help. Almost 1/2 of kfrends return with bad Q=0
camac errors. Also, lots of SDAQ readback errors

Page the run manager.

Eventually, we figured out that crate 23 was going
over-temperature, and power-cycling itself. This
was due to the fan in the power supply dying.

We accessed prep and got a spare and swapped it
out. Cool running now.

14386 23:45 Start new run.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Get flaky power supply from hall, return to prep, do paperwork.

Regard, Jim